This question is not from a text book but from a homework handout. Can somebody help please. I just don't understand it. Here's the question:
Angle PQR and angle ABC are complementary angles and angle PQR is nine times as large as angle ABC. Determine the measure of each angle.

Answer by Mathtut(3670)

complementary angles by definition are equal to ninety degrees
:
lest call angle PQR-P and angle ABC lets call A
:
P+A=90.......eq 1
P=9A.........eq 2
:
plug P's value from eq 2 into eq 1
:
9A+A=90
:
10A=90
:
A=9 degrees
:
P=9A=9(9)=81 degrees